What volume (in mL) of 0.50 M potassium nitrate contains 0.021 mol of potassium nitrate?


1
Expert's answer
2021-04-06T06:07:41-0400

M=mol/volume

Volume=mol/M

V= 0.021÷0.5 = 0.042 L

Volume in Ml = 0.042×1000

Ans= 42mL